OpenStudy (anonymous):

Solve the equation. 4 times the square root of x minus 5 minus 8 = 12 A. 20 B. 30 C. 25 D. 25/4

OpenStudy (anonymous):

4*\[4\times \sqrt{X }-5-8=12?\]

OpenStudy (anonymous):

\[4\sqrt{x-5}-8=12\]

OpenStudy (anonymous):

B